Let X and Y have Joint probability function given by f(x,y) = xy/66 x = 2,3,4 ; y = 1,2,3 Form the p.d table and find marginal probabilities function of X and Y? Also a) Find P (X=4 | Y=2)? b) Find P (X + Y < 3)? c) Are X and Y independent?


1
Expert's answer
2022-02-07T17:59:38-0500

The probability that "X" takes value "x" and "Y" takes value "y" is: "p_{x,y}=f(x,y)=\frac{xy}{66}". The probability distribution table has the form:

"\begin{array}{ c c c }
 (x,y) & p_{x,y}\\
 (2,1) & \frac{2}{66} \\
 (2,2) & \frac{4}{66} \\
(2,3) & \frac{6}{66} \\
(3,1) & \frac{3}{66} \\
 (3,2) & \frac{6}{66} \\
(3,3) & \frac{9}{66} \\
(4,1) & \frac{4}{66} \\
 (4,2) & \frac{8}{66} \\
(4,3) & \frac{12}{66} \\
 \end{array}"

As we can see from the table, "\frac{1}{66}(2+4+6+3+6+9+4+8+12)=\frac{54}{66}" . Since the sum is less than 1, "X", "Y" with the function "f(x,y)" is not a proper distribution. For correct distribution the formulae have the form:

a). "P(X=4|Y=2)=\frac{P((X=4)\cap(Y=2))}{P(Y=2)}=\frac{f(4,2)}{P((X=2)\cap(Y=2))
+P((X=3)\cap(Y=2))+P((X=4)\cap(Y=2))}="

"=\frac{f(4,2)}{f(2,2)
+f(3,2)+f(4,2)}"

b). "P(X+Y<3)=P(X+Y=2)=0" It follows from fact that the minimum value of "X" is 2.

c). By definition, two events "A" and "B" are independent if and only if "P(A\cap B)=P(A)P(B)". It is enough to check that "P(X=x,Y=y)=f(x,y)=P(X=x)P(Y=y)=(\sum_{l}f(x,l))(\sum_{k}f(k,y))"